OpenStudy (anonymous):

What are the next three terms in the sequences below? 2, 5, 7, 12, ___, ___, ___

OpenStudy (anonymous):

Can you determine the type of this sequence first ?

OpenStudy (anonymous):

Subtraction?

OpenStudy (anonymous):

The next number is the sum of the last two. Now do the additions.

OpenStudy (anonymous):

I am so confused

OpenStudy (anonymous):

2+5 = 7 5+7 = 12 7+12 = 19 12 + 19 = ? . . .

OpenStudy (anonymous):

31

OpenStudy (anonymous):

Those are the answers then. 19, 31, 50